linear algebra help.

Let A = 10 1 0 A.) Show s, is LI and forms a basis of R. - Dے واک وه ک ( ه ا ك (.) ( 2/ 5 24 (5) رک) pan 3 | 2 | - |

S2=AV1,AV2,AV3)
Answer #1

FI 0 07 Let A = 1o. l ol Si={ vi, ve, Us} whele Vi=107 V TO Fil V2 = 10 . (A) Let ai, az and q be the three scalalf such that(2) from equations (2) and (3), we have Cla= 0, 0+43 = 0 =9330 from equation (I), we have anto 30 9.30 We have a=0, a zo, a =(3) three linecally independent vectors is a basis of the vector Space R² (B) Sa= favo, Ave, Avs? Now AV, Il lolo Fitoto1 fil(4) Pri] 1907 A ? LJ, LJ, L Let by ba and by be the three scalabs such that bilavi) + ba (Avs) + b3 14V3)=0 J Litir in To 0 =(5) Ź bi = -1 we have bi=-7, ba zo. and b3 = 1 Since call the thifce scalabs bi, ba and b3 cole not equal to zero, the lefole(6) equating the corresponding Co-ordinates We have Git %3=1 417) 92 = -18) att 3 =4 (9) flom (8) and (9), we have alitat24 a(d) Let Y, Y, Y3 be the three scalaus such that TH 21=yout Savat Yz Uz - - O > St Ys7 I ya = |_ Yat 93 equating the correspon.. we have 3 = (-1) Vita) Vg +1) V3 14 Hence the vectol is expressed [4] as a linear Combination of the vectors Vi, Va and Uz